1. Quick Company acquired a piece of equipment in Year 1 at a cost of $100,000. The equipment has a 10-year estimated life, zeroIf Quick Company used Double Declining Balance (DDB) depreciation method instead of straight-line, calculate the following:
    1. Depreciation expense each year
    2. Accumulated depreciation each year
    3. Net book value each year
    4. Impairment loss (if any) at the end of year 4
    5. Comparing the impairment loss in d) with the impairment loss we calculated in class under the straight-line method, discuss the implication.
Answer #1

Answer for a, b, c :

Step 1 : calculation of Depreciation under double declining method

Depreciation expense = Depreciation rate × opening book value of asset

Depreciation rate = 2 × 1/10 × 100 = 20%

Year opening book vale Depreciation rate Depreciation expense Accumulated depreciation closing book value
1 $1,00,000 20% $20,000 $20,000 $80,000
2 $80,000 20% $16,000 $36,000 $64,000
3 $64,000 20% $12,800 $48,800 $51,200
4 $51,200 20% $10,240 $59,040 $40,960

Answee for d :

Step 2 : Calculation of impairment loss

Impairment loss = recoverable Amount - carrying amount

i) Recoverable amount :

Recoverable amount is higher of the following :

1) Fair Value less cost of disposal
2) Value in use (Expected future cash flows at discount rate)

That implies,

1) $50,000

2) $51,000, which ever higher is recoverable Amount

Therefore

Recoverable amount = $51,000

ii) carrying amount under Double declining method = $40,960

Conclusion : there is no impairment loss as recoverable amount is greater than carrying amount under double declining method.

Answer for e :

1) impairment loss under straight line method

i) recoverable amount = $51,000 (same as we calculated above)

ii) carrying amount = cost of asset - Depreciation

Depreciation for 4 years = (cost of asset - salvage value) ÷ life of asset

= ($1,00,000 ÷ 10) × 4 = $40,000

Carrying amount = $1,00,000 -$40,000 = $60,000

Therefore

Impairment loss = recoverable Amount - carrying amount

= $51,000 - $60,000

= $9,000

Conclusion : As we observed that when quick Company used DDB method for Depreciation then there is no impairment loss but when it used straight line method for Depreciation it resulted $9,000 of impairment loss.